The area of a triangle is 8.25 square centimeters. If the base of the triangle is 3 centimeters, what is the height.
Answer by Maths68(1474)

Given
Area=8.25
base=b=3cm
height=h=?
Area of Triangle = (base*hieght)/2
A=(b*h)/2
8.25=(3*h)/2
8.25 * 2=3*h
16.5=3h
or
3h=16.5
h=16.5/3
h=5.5cm